a number exceeds another number by 5.the sum of the number is 19. find the number
Answer by stanbon(48558)

a number exceeds another number by 5.the sum of the number is 19. find the number
--------------
Let one of the numbers be "x"; then the other is "x+5".
Equation:
x + x+5 = 19
2x = 14
x = 7 (the 1st number)
x+5 = 12 (the other number)
================================
Cheers,
Stan H.